It takes a plane 40 min. longer to fly from boston to los angeles at 525 mi/h than it does to return at 600 mi/h. how far apart are the cities?
Answer by Paul(988)

Let x be distance
525(x+40)=600x
525x+21000=600x
525x+21000-525x=600x-525x
21000=75x (divide both sides by 7)
280 = x
Therefore they are 280 miles apart.